The second angle in a triangle is one and a half times the measure of the first. The third is 15 more than half the first angle. Find the measure of each angle in the triangle
Answer by stanbon(75887)

The second angle in a triangle is one and a half times the measure of the first. The third is 15 more than half the first angle. Find the measure of each angle in the triangle
---
1st: x
2nd: (3/2)x
3rd:(x/2)+15
----
Equation:
x + (3/2)x + (x/2)+15 = 180
3x + 15 = 180
3x = 165
x = 55 degrees
---
1st: x = 55 degrees
2nd (3/2)x = 82.5 degrees
3rd: (x/2)+15 = 42.5 degrees
=================================
Cheers,
Stan H.
=========